# J-HERTz: J-PLUS Heritage Exploration of Radio Targets at z 5

### v1

@@ 3.2 Galaxies
We have 390k galaxies in our catalog, 320k with mag. The best redshifts and radio properties are shown in Fig. 6. Around 72% of the best-zs are below 0.5, the lowest belonging to NGC 2976 (a small galaxy in the M81 group) at z = . Only 0.50% are above 1, and many of them are spectroscopically classified as QSOs or have unreliable spec-zs, which can be filtered using the incompatible_spec_class and unreliable_spec_z flags). The radio luminosities peak at , with most falling between and and some lower outliers than can also be filtered with the same flags. The estimated linear size of the galaxies’ radio emission for Resolved sources with an angular size larger than 6 arcsec, mainly falls below 3000 kpc, with the most extended radio source reaching 4.9 Mpc and 12 arcmin (WISEA J003623.89+253611.1), and the second largest with 4.4 Mpc and 40 arcmin being 3C 236, one of the largest known giant radio galaxies (Shulevski2019). Using the same optical resolution criteria as defined in Sect. 3.1 for the hosts, approximately 87% of the galaxies are resolved in J-PLUS, a much higher percentage than the QSOs. This highlights the effectiveness of BANNJOS in using optical morphology to classify sources. Their axis ratio distribution is wider as well, with 40% of sources having an axis ratio , indicating more varied morphological types. Taking only the cases with a relative error below 10% in and bands, galaxies have a median and dispersion () color of , much redder than QSOs, but showing a lot of variation from bluer () up to redder () colors. The radio-loudness in galaxies appears to have a single distribution with a peak around /, but it is skewed towards higher values and reaches /.
@@
The right panel of Fig. 5 shows the distribution of stellar masses of galaxies (Sect. 2.9). We filter for galaxies with mag, masses with low uncertainty and redshift values with no flags. Most galaxies have a stellar mass between and , peaking at . The few outliers, although passing all the filters, were adjusted from the original values from Arizo2025 with a new z_best that differs considerably from the original J-PLUS photo-z used and should not be trusted. For example, the cases with extremely low masses were classified as stars in SDSS and have spectroscopic redshifts compatible with zero, which when used to adjust the original masses resulted in these low values. On the other hand, the sources with high masses were classified as QSOs in SDSS and have much higher spectroscopic redshifts. These problematic cases can be avoided using again the incompatible_spec_class flag. In any case, the fraction of galaxies with stellar masses outside is only 0.11%.
3.3 Stars
@@
Our initial cross-matched sample includes 21.6k objects classified as stars by BANNJOS with associated radio emission. However, most of these are expected to be spurious associations (Callingham2019), and therefore a series of filters are applied to identify a more reliable set of candidates. First, we remove objects with inconsistent spectroscopic classifications, multiple BANNJOS identifications, or with a nearby galaxy or QSO within the radio emission region (see Sect. 4.4). This step eliminates most extended radio sources incorrectly associated with stars. After filtering, we retain 9k extended radio sources with Resolved = False and only 347 Resolved stars with extended radio emission, likely contaminated by nearby undetected sources. We also recover 5k compact (, Resolved=False) sources. This last group constitutes the most promising candidate radio stars. Because genuine radio stars are intrinsically faint (typical – W Hz-1; Callingham2021; Driessen2024), LoTSS flux limits constrain the distance where detection is possible to 1 kpc of the Sun. We cross-match our compact, filtered stars with Gaia DR3, and retain only sources with reliable parallaxes () and distances kpc, using the catalog in BailerJones2021. This yields a final sample of 831 stars, with inferred radio luminosities consistent with the expected range. To assess the possibility of chance alignments between radio AGN and optical stars, we follow the likelihood-ratio method of Callingham2019 and Sutherland1992. Compared with offset positions, we find an excess of cross-matches within 5′′ by a factor 25–40, indicating that the majority of these associations are unlikely to be spurious. The radio-loudness distribution of our candidates is significantly narrower and more radio-quiet than that of the full star sample, consistent with expectations for nearby stars (left panel of Fig. 7). In the Gaia color–magnitude diagram (right panel of Fig. 7), our candidates populate the main sequence and display a luminosity gradient similar to that reported for confirmed radio stars in the Sydney Radio Star Catalogue (Fig. 4 in Driessen2024). Since that catalogue is dominated by southern-hemisphere objects, only two sources are cross-matched with our sample (one of which is highlighted in Fig. 8). Both stars are also included in the sample of M dwarfs observed at low radio frequencies by Callingham2021. Although the overlap is small, this agreement provides encouraging support for the reliability of our candidate selection. We therefore propose this final set of 831 sources as candidate radio stars. These are marked with a radio_star_candidate = ‘reliable’ in the final catalog, while the rest of the stars have an ‘unreliable’ value. A more detailed characterization is required, however, before any of them can be confirmed as genuine detections. Possible mechanisms to explain the radio continuum emission in these sources include stellar coronal activity (Callingham2021) or magnetic interactions with close-in exoplanets (Pope2021), although further follow-up will be needed to establish the nature of these objects.
4 Discussion
@@
In this section we discuss the main properties of galaxies and QSOs, focusing on their radio and IR continuum. Specifically, the results based on improved photo-zs and stellar masses derived with narrow-band photometry are compared with those obtained by Hardcastle2023; Hardcastle2025 for the larger LoTSS DR2 sample. Additionally, we analyze the radio-loudness properties for the different populations identified by BANNJOS. In the case of galaxies, we discuss how the radio activity relates with the masses and SFRs. Finally, in Section 4.4 we discuss some caveats and user warnings for the catalog.
4.1 Radio luminosities, loudness and sizes
@@
The lower left panel in Fig. 6 shows that QSOs radio luminosities peak at a higher value when compared to galaxies, possibly due to a selection effect given the different redshift distributions of the two populations. Limiting the comparison only to QSOs with z_best 1, they peak at a much closer value of , although this remains slightly higher than for galaxies. Only 2% of the galaxies would fall within the Faranoff-Riley II (FRII) luminosity cut (; FaranoffRiley1974), while that percentage is 40% for QSOs. The left panel of Fig. 9 illustrates the relationship between radio luminosity and estimated radio linear sizes for Resolved sources with an angular size larger than 6 arcsec. Many of the galaxies fall in the Faranoff-Riley I (FRI) region, where the radio emission linear size is strongly correlated with the radio luminosity. On the other hand, most QSOs fall in the FRII region and their radio size and luminosity are not strongly correlated. All sources reach more extended radio emission in the FRII region, typically due to larger jets and brighter radio lobes at their edges, but this FRI/FRII threshold does not appear to divide the sample in a meaningful way. Moreover, the lack of low-luminosity extended sources is most likely caused by selection effects, as shown by the solid blue line indicating the surface brightness limit of LoTSS (see fig. 8 in Hardcastle2025). The distribution of sizes and luminosities in our sample is consistent with the radio-luminous AGN population in their work.
@@
The lower right panel of Fig. 6 shows that QSOs seem to exhibit a bimodal distribution in radio-to-IR radio-loudness. The majority of sources peak at lower ratios, with an extended tail or possibly a second distribution at higher ratios. This bimodality has been investigated in previous works using radio-to-optical ratios or radio luminosities as radio-loudness indicators, with mixed results. Kellermann1989; Laor2000; Ivezic2002; Kellermann2016; Balokovic2012; Zhang2021 and Arnaudova2024 find that the radio-loudness distribution can be best fitted with two Gaussian components, or that there are different properties among the two populations, even if the separation is not a clear division. On the other hand, White2000 and Cirasuolo2003 find continuous distributions from low to high radio-loudness without a clear gap, but their analysis is limited to samples of a few hundred QSOs. Mahony2012 find a bimodal distribution for a sample covering a wide range in redshift, which is interpreted as evolutionary effect rather than an intrinsic property of the QSO population.
@@
We perform an analysis using the Bayesian Information Criterion (BIC) and the Akaike Information Criterion (AIC), comparing a 2-normal distribution with different continuous asymmetric distributions such as the log-normal, skewed-normal and exponential-normal. To avoid possible selection effects, we limit the sample to QSOs with z_best 1. The 2-normal distribution is favored over the others by both diagnostics, with and 500. The best fit for the whole QSO sample provides a Gaussian distribution centered at with , and a second one centered at with . The radio-loudness value at the intersection of these Gaussians is adopted as the threshold between both populations, i.e. at /. We consider sources below this threshold as radio-quiet, and radio-loud above it. However, we note that this is not an absolute division, as the tails of both populations overlap. 5% of QSOs are radio-loud at low redshifts (z 0.2) and the fraction increases to 12% at z 1, which is very similar to the usual fraction of radio-loud QSOs observed at low redshifts in the literature (Kellermann1989; Ho2002). The radio-loud fraction reaches a maximum of 20% at z = 3.5 – 4.0, but selection effects could come into play at higher redshifts. Moreover, the exact fraction of radio-loud objects will depend on the chosen definition for radio-loudness and, as mentioned before, there is still a debate about whether there is a real bimodality in the radio-loudness distribution of QSOs.
@@
Galaxies in our sample show radio-loudness values centered around , without exhibiting the bimodality observed in QSOs. Overall, only 5% of galaxies fall into the radio-loud category. Their high radio-to-infrared flux ratios suggest that these nuclei have undergone intense jet activity in the past, consistent with the long timescales traced by low-frequency synchrotron emission. Nevertheless, these galaxies show no clear signatures of ongoing nuclear activity in their J-PLUS photometry or WISE colors. To study them in more detail, we cross-match our radio-loud galaxies with the catalog by Drake2024, who provide LoTSS sources with SDSS emission line spectra from the Portsmouth catalog (Thomas2013), radio-excess probabilities, and spectral classifications of a subset of hosts with full line diagnostics in a BPT diagram (Baldwin1981). J-HERTz has 3,621 radio-loud galaxies in common with their sample. Following the classification by Drake2024, galaxies are identified as Low-Excitation Radio Galaxies (LERGs; Laing1994; Heckman2014) if they exhibit radio-excess and fall within the LINER, composite, or star-forming regions in the BPT diagram. The sources located below the Kewley2001 star-forming boundary can be considered as optically quiescent LERGs, i.e. radio-loud galaxies with a weak or negligible contribution from the AGN to the nebular ionized gas. Among the common sample, 12% of radio-loud galaxies are LERGs falling below the star-forming boundary, 5% are compatible with HERGs, and 45% lack a clear classification or miss any of the BPT lines (H, H, [O III] or [N II]).
@@
Nevertheless, the total fraction of optically quiescent LERGs should also include passive galaxies without significant line emission that cannot be classified in the BPT diagram, such as optically inactive radio galaxies. Following the criteria used in Stasinska2025, a host is considered inactive if its H equivalent width is less than 3Å. Among the radio-loud galaxies without a BPT classification but with measured H equivalent widths (and continuum S/N 3), 7% are found to be optically inactive. This implies that the total fraction of optically quiescent LERGs is approximately 20% of the radio-loud galaxy population. It is interesting to compare this result with the fraction of remnant AGN, i.e. sources in which jet activity at the nucleus has completely ceased, usually identified as radio galaxies with undetected radio cores in deep surveys or with ultra-steep spectral indices. Previous studies have reported upper limits of 5–10% for the fraction of AGN remnants among radio-loud galaxies (Mahatma2018; Dutta2023; Mostert2023). This suggests that most optically quiescent LERGs still host an active radio core that is unable to ionize the gas but continues to power the radio lobes. Identifying these systems is crucial for understanding the AGN duty cycle, as they represent a substantial fraction of the radio-loud population (20%) showing negligible or undetectable signs of nuclear activity in the optical, yet with ongoing jet activity that may contribute significantly to the black hole accretion history.
@@ 4.2 IR properties
The IR emission provides a powerful diagnostic to distinguish star-forming galaxies from radio-loud AGN. In the radio–IR plane (left panel of Fig. 10), most galaxies follow the well-known IR–radio relation (Helou1985), consistent with radio emission from star formation activity (see Fernández Gil et al., in prep., for a discussion of the SFR-radio relation in J-HERTz galaxies). QSOs, on the other hand, occupy the brightest IR region, characterized by strong dust emission. Among them, only a fraction are radio loud, showing radio excess above the IR–radio relation () together with radio-loud galaxies, while most QSOs are radio quiet and overlap in luminosity with the star-forming galaxy population.
@@
The WISE color–color diagram (Fig. 10, right) further distinguishes these populations. The W1-W2 color traces hot dust heated by AGN activity, while W2-W3 traces warm dust from star-formation. Passive galaxies occupy the bluer region of the diagram, as they show little IR emission in either color, whereas star-forming galaxies show redder W2–W3 values, and AGN/QSOs have the redder W1–W2 colors. One of the main parameters used by BANNJOS to identify QSOs is indeed W1-W2 (e.g. Fig. 11(a) for a QSO with red W1-W2 color; Fig. 11(b) for a galaxy with blue W1-W2 color). A similar WISE color–color diagram is shown in Hardcastle2025, where star-forming galaxies, radio-excess AGN, and spectroscopically confirmed QSOs occupy regions consistent with those found in our sample, confirming that mid-IR colors are effective in distinguish between different source types.
@@
The radio-loudness versus W1–W2 color plane (Fig. 12) highlights these differences. It is divided in 4 regions by two dashed lines: the radio-loud/radio-quiet threshold defined earlier at /, and a blue/red division at W1-W2 = , obtained by fitting two components to the radio-loud galaxies and QSOs, which are separated by their W1-W2 colors. Radio-loud galaxies exhibit relatively blue W1–W2 colors, consistent with LERGs or optically inactive radio galaxies, disfavoring scenarios involving dusty, obscured Type 2 QSOs (Gurkan2014; Mingo2016). QSOs populate mainly the reddest colors. Most radio-quiet galaxies have blue W1-W2 colors, but a subset shows higher IR colors comparable to those of QSOs. Approximately 48% of the galaxies with W1-W2 0.25 and available spectra are classified as QSOs in SDSS, suggesting that many of them are likely identified as galaxies due to their resolved or extended morphologies. Alternatively, some could be starburst galaxies with intense star formation heating their dust, or systems dominated by older stellar populations that emit strongly in the near-IR. These scenarios will be explored further in the next section.
4.3 Stellar masses and star formation rates of galaxies

### v2

@@ III.2 Galaxies
We have 390k galaxies in our catalog, 320k with mag. The best redshifts and radio properties are shown in Fig. 6. Around 72% of the best-zs are below 0.5, the lowest belonging to NGC 2976 (a small galaxy in the M81 group) at z = . Only 0.50% are above 1, and many of them are spectroscopically classified as QSOs or have unreliable spec-zs, which can be filtered using the incompatible_spec_class and unreliable_spec_z flags). The radio luminosities peak at , with most falling between and and some lower outliers than can also be filtered with the same flags. The estimated linear size of the galaxies’ radio emission for Resolved sources with an angular size larger than 6 arcsec, mainly falls below 3000 kpc, with the most extended radio source reaching 4.9 Mpc and 12 arcmin (WISEA J003623.89+253611.1), and the second largest with 4.4 Mpc and 40 arcmin being 3C 236, one of the largest known giant radio galaxies (Shulevski et al., 2019). Using the same optical resolution criteria as defined in Sect. III.1 for the hosts, approximately 87% of the galaxies are resolved in J-PLUS, a much higher percentage than the QSOs. This highlights the effectiveness of BANNJOS in using optical morphology to classify sources. Their axis ratio distribution is wider as well, with 40% of sources having an axis ratio , indicating more varied morphological types. Taking only the cases with a relative error below 10% in and bands, galaxies have a median and dispersion () color of , much redder than QSOs, but showing a lot of variation from bluer () up to redder () colors. The radio-loudness in galaxies appears to have a single distribution with a peak around /, but it is skewed towards higher values and reaches /.
@@
The right panel of Fig. 5 shows the distribution of stellar masses of galaxies (Sect. II.9). We filter for galaxies with mag, masses with low uncertainty and redshift values with no flags. Most galaxies have a stellar mass between and , peaking at . The few outliers, although passing all the filters, were adjusted from the original values from Arizo-Borillo et al. (2025) with a new z_best that differs considerably from the original J-PLUS photo-z used and should not be trusted. For example, the cases with extremely low masses were classified as stars in SDSS and have spectroscopic redshifts compatible with zero, which when used to adjust the original masses resulted in these low values. On the other hand, the sources with high masses were classified as QSOs in SDSS and have much higher spectroscopic redshifts. These problematic cases can be avoided using again the incompatible_spec_class flag. In any case, the fraction of galaxies with stellar masses outside is only 0.11%.
III.3 Stars
@@
Our initial cross-matched sample includes 21.6k objects classified as stars by BANNJOS with associated radio emission. However, most of these are expected to be spurious associations (Callingham et al. 2019), and therefore a series of filters are applied to identify a more reliable set of candidates. First, we remove objects with inconsistent spectroscopic classifications, multiple BANNJOS identifications, or with a nearby galaxy or QSO within the radio emission region (see Sect. IV.4). This step eliminates most extended radio sources incorrectly associated with stars. After filtering, we retain 9k extended radio sources with Resolved = False and only 347 Resolved stars with extended radio emission, likely contaminated by nearby undetected sources. We also recover 5k compact (, Resolved=False) sources. This last group constitutes the most promising candidate radio stars. Because genuine radio stars are intrinsically faint (typical – W Hz-1; Callingham et al. 2021; Driessen et al. 2024), LoTSS flux limits constrain the distance where detection is possible to 1 kpc of the Sun. We cross-match our compact, filtered stars with Gaia DR3, and retain only sources with reliable parallaxes () and distances kpc, using the catalog in Bailer-Jones et al. (2021). This yields a final sample of 831 stars, with inferred radio luminosities consistent with the expected range. To assess the possibility of chance alignments between radio AGN and optical stars, we follow the likelihood-ratio method of Callingham et al. (2019) and Sutherland and Saunders (1992). Compared with offset positions, we find an excess of cross-matches within 5′′ by a factor 25–40, indicating that the majority of these associations are unlikely to be spurious. The radio-loudness distribution of our candidates is significantly narrower and more radio-quiet than that of the full star sample, consistent with expectations for nearby stars (left panel of Fig. 7). In the Gaia color–magnitude diagram (right panel of Fig. 7), our candidates populate the main sequence and display a luminosity gradient similar to that reported for confirmed radio stars in the Sydney Radio Star Catalogue (Fig. 4 in Driessen et al., 2024). Since that catalogue is dominated by southern-hemisphere objects, only two sources are cross-matched with our sample (one of which is highlighted in Fig. 8). Both stars are also included in the sample of M dwarfs observed at low radio frequencies by Callingham et al. (2021). Although the overlap is small, this agreement provides encouraging support for the reliability of our candidate selection. We therefore propose this final set of 831 sources as candidate radio stars. These are marked with a radio_star_candidate = ‘reliable’ in the final catalog, while the rest of the stars have an ‘unreliable’ value. A more detailed characterization is required, however, before any of them can be confirmed as genuine detections. Possible mechanisms to explain the radio continuum emission in these sources include stellar coronal activity (Callingham et al., 2021) or magnetic interactions with close-in exoplanets (Pope et al., 2021), although further follow-up will be needed to establish the nature of these objects.
IV Discussion
@@
In this section we discuss the main properties of galaxies and QSOs, focusing on their radio and IR continuum. Specifically, the results based on improved photo-zs and stellar masses derived with narrow-band photometry are compared with those obtained by Hardcastle et al. (2023, 2025) for the larger LoTSS DR2 sample. Additionally, we analyze the radio-loudness properties for the different populations identified by BANNJOS. In the case of galaxies, we discuss how the radio activity relates with the masses and SFRs. Finally, in Section IV.4 we discuss some caveats and user warnings for the catalog.
IV.1 Radio luminosities, loudness and sizes
@@
The lower left panel in Fig. 6 shows that QSOs radio luminosities peak at a higher value when compared to galaxies, possibly due to a selection effect given the different redshift distributions of the two populations. Limiting the comparison only to QSOs with z_best 1, they peak at a much closer value of , although this remains slightly higher than for galaxies. Only 2% of the galaxies would fall within the Faranoff-Riley II (FRII) luminosity cut (; Fanaroff and Riley, 1974), while that percentage is 40% for QSOs. The left panel of Fig. 9 illustrates the relationship between radio luminosity and estimated radio linear sizes for Resolved sources with an angular size larger than 6 arcsec. Many of the galaxies fall in the Faranoff-Riley I (FRI) region, where the radio emission linear size is strongly correlated with the radio luminosity. On the other hand, most QSOs fall in the FRII region and their radio size and luminosity are not strongly correlated. All sources reach more extended radio emission in the FRII region, typically due to larger jets and brighter radio lobes at their edges, but this FRI/FRII threshold does not appear to divide the sample in a meaningful way. Moreover, the lack of low-luminosity extended sources is most likely caused by selection effects, as shown by the solid blue line indicating the surface brightness limit of LoTSS (see fig. 8 in Hardcastle et al. 2025). The distribution of sizes and luminosities in our sample is consistent with the radio-luminous AGN population in their work.
@@
The lower right panel of Fig. 6 shows that QSOs seem to exhibit a bimodal distribution in radio-to-IR radio-loudness. The majority of sources peak at lower ratios, with an extended tail or possibly a second distribution at higher ratios. This bimodality has been investigated in previous works using radio-to-optical ratios or radio luminosities as radio-loudness indicators, with mixed results. Kellermann et al. (1989); Laor (2000); Ivezić et al. (2002); Kellermann et al. (2016); Baloković et al. (2012); Zhang et al. (2021) and Arnaudova et al. (2024) find that the radio-loudness distribution can be best fitted with two Gaussian components, or that there are different properties among the two populations, even if the separation is not a clear division. On the other hand, White et al. (2000) and Cirasuolo et al. (2003) find continuous distributions from low to high radio-loudness without a clear gap, but their analysis is limited to samples of a few hundred QSOs. Mahony et al. (2012) find a bimodal distribution for a sample covering a wide range in redshift, which is interpreted as evolutionary effect rather than an intrinsic property of the QSO population.
@@
We perform an analysis using the Bayesian Information Criterion (BIC) and the Akaike Information Criterion (AIC), comparing a 2-normal distribution with different continuous asymmetric distributions such as the log-normal, skewed-normal and exponential-normal. To avoid possible selection effects, we limit the sample to QSOs with z_best 1. The 2-normal distribution is favored over the others by both diagnostics, with and 500. The best fit for the whole QSO sample provides a Gaussian distribution centered at with , and a second one centered at with . The radio-loudness value at the intersection of these Gaussians is adopted as the threshold between both populations, i.e. at /. We consider sources below this threshold as radio-quiet, and radio-loud above it. However, we note that this is not an absolute division, as the tails of both populations overlap. 5% of QSOs are radio-loud at low redshifts (z 0.2) and the fraction increases to 12% at z 1, which is very similar to the usual fraction of radio-loud QSOs observed at low redshifts in the literature (Kellermann et al., 1989; Ho, 2002). The radio-loud fraction reaches a maximum of 20% at z = 3.5 – 4.0, but selection effects could come into play at higher redshifts. Moreover, the exact fraction of radio-loud objects will depend on the chosen definition for radio-loudness and, as mentioned before, there is still a debate about whether there is a real bimodality in the radio-loudness distribution of QSOs.
@@
Galaxies in our sample show radio-loudness values centered around , without exhibiting the bimodality observed in QSOs. Overall, only 5% of galaxies fall into the radio-loud category. Their high radio-to-infrared flux ratios suggest that these nuclei have undergone intense jet activity in the past, consistent with the long timescales traced by low-frequency synchrotron emission. Nevertheless, these galaxies show no clear signatures of ongoing nuclear activity in their J-PLUS photometry or WISE colors. To study them in more detail, we cross-match our radio-loud galaxies with the catalog by Drake et al. (2024), who provide LoTSS sources with SDSS emission line spectra from the Portsmouth catalog (Thomas et al., 2013), radio-excess probabilities, and spectral classifications of a subset of hosts with full line diagnostics in a BPT diagram (Baldwin et al., 1981). J-HERTz has 3,621 radio-loud galaxies in common with their sample. Following the classification by Drake et al. (2024), galaxies are identified as Low-Excitation Radio Galaxies (LERGs; Laing et al., 1994; Heckman and Best, 2014) if they exhibit radio-excess and fall within the LINER, composite, or star-forming regions in the BPT diagram. The sources located below the Kewley et al. (2001) star-forming boundary can be considered as optically quiescent LERGs, i.e. radio-loud galaxies with a weak or negligible contribution from the AGN to the nebular ionized gas. Among the common sample, 12% of radio-loud galaxies are LERGs falling below the star-forming boundary, 5% are compatible with HERGs, and 45% lack a clear classification or miss any of the BPT lines (H, H, [O III] or [N II]).
@@
Nevertheless, the total fraction of optically quiescent LERGs should also include passive galaxies without significant line emission that cannot be classified in the BPT diagram, such as optically inactive radio galaxies. Following the criteria used in Stasińska et al. (2025), a host is considered inactive if its H equivalent width is less than 3Å. Among the radio-loud galaxies without a BPT classification but with measured H equivalent widths (and continuum S/N 3), 7% are found to be optically inactive. This implies that the total fraction of optically quiescent LERGs is approximately 20% of the radio-loud galaxy population. It is interesting to compare this result with the fraction of remnant AGN, i.e. sources in which jet activity at the nucleus has completely ceased, usually identified as radio galaxies with undetected radio cores in deep surveys or with ultra-steep spectral indices. Previous studies have reported upper limits of 5–10% for the fraction of AGN remnants among radio-loud galaxies (Mahatma et al., 2018; Dutta et al., 2023; Mostert et al., 2023). This suggests that most optically quiescent LERGs still host an active radio core that is unable to ionize the gas but continues to power the radio lobes. Identifying these systems is crucial for understanding the AGN duty cycle, as they represent a substantial fraction of the radio-loud population (20%) showing negligible or undetectable signs of nuclear activity in the optical, yet with ongoing jet activity that may contribute significantly to the black hole accretion history.
@@ IV.2 IR properties
The IR emission provides a powerful diagnostic to distinguish star-forming galaxies from radio-loud AGN. In the radio–IR plane (left panel of Fig. 10), most galaxies follow the well-known IR–radio relation (Helou et al., 1985), consistent with radio emission from star formation activity (see Fernández Gil et al., in prep., for a discussion of the SFR-radio relation in J-HERTz galaxies). QSOs, on the other hand, occupy the brightest IR region, characterized by strong dust emission. Among them, only a fraction are radio loud, showing radio excess above the IR–radio relation () together with radio-loud galaxies, while most QSOs are radio quiet and overlap in luminosity with the star-forming galaxy population.
@@
The WISE color–color diagram (Fig. 10, right) further distinguishes these populations. The W1-W2 color traces hot dust heated by AGN activity, while W2-W3 traces warm dust from star-formation. Passive galaxies occupy the bluer region of the diagram, as they show little IR emission in either color, whereas star-forming galaxies show redder W2–W3 values, and AGN/QSOs have the redder W1–W2 colors. One of the main parameters used by BANNJOS to identify QSOs is indeed W1-W2 (e.g. Fig. 11(a) for a QSO with red W1-W2 color; Fig. 11(b) for a galaxy with blue W1-W2 color). A similar WISE color–color diagram is shown in Hardcastle et al. (2025), where star-forming galaxies, radio-excess AGN, and spectroscopically confirmed QSOs occupy regions consistent with those found in our sample, confirming that mid-IR colors are effective in distinguish between different source types.
@@
The radio-loudness versus W1–W2 color plane (Fig. 12) highlights these differences. It is divided in 4 regions by two dashed lines: the radio-loud/radio-quiet threshold defined earlier at /, and a blue/red division at W1-W2 = , obtained by fitting two components to the radio-loud galaxies and QSOs, which are separated by their W1-W2 colors. Radio-loud galaxies exhibit relatively blue W1–W2 colors, consistent with LERGs or optically inactive radio galaxies, disfavoring scenarios involving dusty, obscured Type 2 QSOs (Gürkan et al., 2014; Mingo et al., 2016). QSOs populate mainly the reddest colors. Most radio-quiet galaxies have blue W1-W2 colors, but a subset shows higher IR colors comparable to those of QSOs. Approximately 48% of the galaxies with W1-W2 0.25 and available spectra are classified as QSOs in SDSS, suggesting that many of them are likely identified as galaxies due to their resolved or extended morphologies. Alternatively, some could be starburst galaxies with intense star formation heating their dust, or systems dominated by older stellar populations that emit strongly in the near-IR. These scenarios will be explored further in the next section.
IV.3 Stellar masses and star formation rates of galaxies